Determine the volume of ethyl alcohol required to prepare 250 mL of a 45% by volume solution.

1
Expert's answer
2022-02-28T11:04:10-0500

Let the volume of alcohol taken = x mL.


Volume percent = Volume of alcohol/ volume of solution* 100.


Or 45= "\frac{x}{250}" or X (i.e. volume of ethanol mixed).


= "\frac{250}{100}\times 45"


= 112.5 mL.